OpenStudy (anonymous):

what is the area of a circle whose diameter equals 8m? a.) 64 π mm² b.) 8 π mm² c .) 16 π mm² d.) 32 π mm² e.) none of the above

OpenStudy (anonymous):

Area of circle = πr² Diameter = 2r

OpenStudy (anonymous):

8= 2r r=4 Now plug into the formula: A= πr² A= π(4)² A= 16π Answer: c .) 16 π mm²

OpenStudy (anonymous):

excuse me , how did you get 16 ? ;o

OpenStudy (anonymous):

(4)² = 4*4 = 16

OpenStudy (anonymous):

i get it now , thanks c: